In the past the standard deviation of weights of certain 40.0 oz filled by a machine was 0.25 oz. A random sample of 20 packages showed a standard deviation of 0.32 oz. Test the hypothesis is the apparent increase in variability significant at the 0.05 level of significance
1
Expert's answer
2021-08-03T12:37:47-0400

"H_0:\sigma=0.25."

"H_a:\sigma>0.25."

Test statistic: "\chi^2=\frac{(20-1)0.32}{0.25}=24.32."

Degrees of freedom: "df=20-1=19."

P-value: "p=P(\chi^2>24.32)=0.1842."

Since the p-value is greater than 0.05, fail to reject the null hypothesis.

There is no apparent increase in variability significant at the 0.05 level of significance.